8) Find L10 for f(x) = 22 +3.c +1 from 2 = 1 to 2 = 2.
Answer #1

[ (22 + 3x + 1) da with n = 10 rectangles, using left endpoints. The left Riemann sum (also known as the left endpoint appro19 2 = b 10 Now, we just evaluate the function at the left endpoints: f (x0) = f(a) = f(1) = 5 = 5 f (x1) = f 21) == ( ) 551241 - - 9.64 5 25 (73) =s () (1) f (29) = f = 1031 100 - = 10.31 Finally, just sum up the above values and 1 multiply by Δα 1

Therefore, the value of L10 is 7.535

ALL THE BEST...!!